# Pedestrian Tracking with Monocular Camera using Unconstrained 3D Motion Model

### v1

@@ Abstract
A first-principle single-object model is proposed for pedestrian tracking. It is assumed that the extent of the moving object can be described via known statistics in 3D, such as pedestrian height. The proposed model thus need not constrain the object motion in 3D to a common ground plane, which is usual in 3D visual tracking applications. A nonlinear filter for this model is implemented using the unscented Kalman filter (UKF) and tested using the publicly available MOT-17 dataset. The proposed solution yields promising results in 3D while maintaining perfect results when projected into the 2D image. Moreover, the estimation error covariance matches the true one. Unlike conventional methods, the introduced model parameters have convenient meaning and can readily be adjusted for a problem.
@@ 1 Introduction
This paper develops a linear object stochastic dynamic model and the corresponding nonlinear measurement model suitable for the monocular pedestrian VTS, that (i) allows unconstrained object motion description in 3D space and (ii) does not rely on the common ground plane assumption. Instead, the proposed BB dynamics model is designed under the assumption of a pedestrian with average body proportions. This idea shares some similarities with [8, 14, 15]. In addition, techniques for setting and identifying the resulting state-space model parameters are discussed and illustrated using available datasets. A numerically stable unscented Kalman filter (UKF) is designed for the proposed model and is analyzed using the MOT-17 dataset [16, 17].
The rest of the paper is organized as follows. Section II provides background for monocular VTSs. Traditional tracking algorithms working in the 2D space are reviewed in Section III. The proposed method for tracking pedestrians in 3D is developed in Section IV. A numerical comparison of the traditional and proposed methods is given in Section V, and concluding remarks are drawn in Section VI.
2 Formalization of the Problem
This section outlines the relation between a moving object in 3D space and its observation via a VDN applied to a camera image. First, the motion in 3D is discussed, followed by the projection of the object at a particular time instant into a pinhole camera image space. The relation of the projection to the output of a VDN is described last.
2.1 Object Motion in 3D
@@
2.2 Pinhole Camera Model
@@
2.2.1 Projection of a 3D Point into Focal Plane
@@
2.2.2 Basis Change to Image Plane
@@
where $\mathbf{r}_{\mathrm{I},\mathrm{F}}$ expressed in pixels is the vector pointing from the origin of the image plane $\mathrm{I}$ to the center of the projection $\mathrm{F}$ as illustrated in Fig. 1b. The velocity $\dot{\mathbf{r}}_{\mathrm{F},\mathrm{P}}$, the length $s_{\mathrm{F}}$ and its temporal change rate $\dot{s}_{\mathrm{F}}$ are transformed from $\mathrm{F}$ to $\mathrm{I}$ by, respectively
@@
2.3 Object’s Bounding Box and Annotation
@@
2.4 Bounding Box Detection Modeling
@@
3 Baseline Filtering Techniques in 2D
@@
3.1 2D-Model-Based Linear Time-Invariant Filter
Papers [5] and [6] present linear time-invariant motion222 The specific model used in this paper is referred to as BB dynamic model 2 in [5], and the notation used herein is slightly varied. and measurement models suitable for VTSs, including identification of their parameters. The motion model from [5] uses the nearly constant velocity (NCV) model, which is described first. Using it as a building block, the state-space model is introduced, followed by the construction of a filter for this model.
3.1.1 Nearly Constant Velocity Model
@@
[…]
3.1.2 State-Space Modeling Based on [5, 6]
@@
$\mathbf{Q}^{\mathrm{2D}}=\text{cov}\!\left(\,\mathbf{w}_{k}^{\mathrm{2D}}\,\right)=\text{diag}\left(\,{q_{\dot{\mathsf{x}}}\mathbf{T},q_{\dot{\mathsf{y}}}\mathbf{T},q_{\dot{\upomega}}\mathbf{T},q_{\dot{\mathsf{h}}}\mathbf{T}}\,\right),$ (31)
@@
3.1.3 2D-Model-Based Filter for (3.1.2)
An optimal unbiased state estimator having minimal mean squared error (MSE) for the linear system such as (3.1.2) can be derived in a closed form. It is the Kalman filter (KF), formed by a recursion composed of the prediction and filtering steps.
@@
where $\mathbf{K}_{k}^{\mathrm{2D}}$ is the Kalman gain. The error corresponding to the predictive and filtering estimates $\mathbf{x}_{k|k\!-\!1}^{\mathrm{2D}}$ and $\mathbf{x}_{k|k}^{\mathrm{2D}}$, respectively, are characterized by the predictive and filtering covariance matrices $\mathbf{P}_{k|k\!-\!1}^{\mathrm{2D}}$ and $\mathbf{P}_{k|k}^{\mathrm{2D}}$, respectively. The filtering step (3.1.3) is performed only when a measurement $\mathbf{z}_{k}$ exists at the time step $k$, which generally applies throughout the paper.
3.1.4 Filter Initialization
[…]
@@ 3.1.4 Filter Initialization
  • The depth can be estimated777 The dept is estimated using (8) with $s_{\mathrm{O}}$ being the pedestrian height, substituted with the assumed mean pedestrian height $m_{h}$. Using (10), $s_{\mathrm{P}}$ is then replaced with the measured height. Note that the height is preferable to the width for the depth estimation, as the width of a pedestrian, including its hands, may change significantly due to hand swinging. as $z_{0|0}=\tfrac{f\cdot m_{h}}{\texttt{|px|}\cdot[\mathbf{z}_{0}]_{4}}$.
@@
3.2 Bag-of-Tricks-Based Filter
@@
A notable adjustment was introduced in DeepSORT [21], which suggested that the covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, appearing in the Kalman filter should be functions of some estimated state or measurement elements (see their code). Presumably, this should mimic the perspective projection effects. The resulting filter, however, is no longer a Kalman filter. The final adjustments introduced in BoT-SORT include changing the original state vector888 The original presentation from BoT-SORT paper [4] and the corresponding code was revisited, leading to alternative notation in this paper. to be
@@
thus all units of the state variables in $\mathbf{x}_{k}^{\mathrm{BoT}}$ (3.2) are pixels.
@@
$\mathbf{F}^{\mathrm{BoT}}$ $=\text{diag}\left(\,{\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right]}\,\right).$ (56)
@@
Notice that the 2D-Model-Based Filter 3.1 completely neglects any 3D modeling, while the Bag-of-Tricks-Based Filter 3.2 takes it heuristically into account up to some level. None of these filters produce estimates in 3D. These issues are tackled in the following Section by modeling directly in 3D.
4 Proposed Model-Based Nonlinear Filter in 3D
@@ 4.1 Geometric Modeling in 3D
The planar 3D BB is a perfectly aligned rectangle with the focal plane, as shown in Fig. 2. All the vertices and edges of the planar 3D BB have the same depth in front of the camera. normal-$\square$
@@
where the position $x_{k}$, $y_{k}$, $z_{k}$ of the lower-bottom center of the box and its width $\omega_{k}$ and height $h_{k}$ are in meters, while the velocities $\dot{x}_{k}$, $\dot{y}_{k}$, $\dot{z}_{k}$ are in meters per second. An example of such a 3D BB state is given in Fig. 2, in which case position is $\mathbf{r}_{\mathrm{C|O}}=[x_{k}\ y_{k}\ z_{k}]^{\top}$, the velocity is $\dot{\mathbf{r}}_{\mathrm{C|O}}=[\dot{x}_{k}\ \dot{y}_{k}\ \dot{z}_{k}]^{\top}$, the height is $s_{\mathrm{O}}=h_{k}$, and the width is not denoted in Fig. 2.
The dynamics (1) of the 3D state is approximated as follows. In accordance with standard motion modeling of physical objects [1], an NCV model from Section 3.1.1 is used to describe the position/velocity of the planar BB in 3D. However, unlike the 2D case, the NCV model does not seem suitable for describing the extent. Therefore, this paper proposes to model the width and height using the following model.
4.1.1 Auto-Regressive Model
@@
where $\tau>0$ is a time-constant101010 $\tau$ can be specified as the elapsed time during which the process $\mathtt{p}(t)$, starting at $\mathtt{p}(0)$, changes its value to $e^{-1}\mathtt{p}(0)+m\approx 0.4\mathtt{p}(0)+m$. , $w(t)$ is a zero-mean continuous white noise, and $\delta(t)$ is the Dirac delta function111111 Similarly to (13), a correct mathematical notation of (60) would involve a stochastic differential equation and a Wiener process [22]. .
@@
where $\alpha=e^{-\nicefrac{{T}}{{\tau}}}$. The above discrete model is usually called the auto-regressive (AR) model of order 1.
4.1.2 State-Space Modeling in 3D
@@
respectively, where $\mathbf{0}_{m\times n}$ is a zero matrix of dimensions $m\times n$, and where $\alpha_{\omega}=e^{-\nicefrac{{T}}{{\tau_{\omega}}}}$ and $\alpha_{h}=e^{-\nicefrac{{T}}{{\tau_{h}}}}$. The parameters of the width and height AR models are as follows. The height of the planar 3D BB is modeled as that of a pedestrian, using the mean $m_{h}=1.65$ m and time-constant $\tau_{h}\approx 4$ s. The width of the planar 3D BB is modeled with the mean $m_{\omega}=0.85$ m, and time-constant $\tau_{\omega}\approx 0.4$ s to account for rough changes due to the hand swinging. The matrix $\mathbf{Q}^{\mathrm{3D}}=\text{cov}\!\left(\,\!\mathbf{w}_{k}^{\mathrm{3D}}\!\,\right)$ is
@@
[…]
The nonlinear measurement equation (62b) is described by the function $\mathbf{x}_{k}^{\mathrm{2D}}=\mathbf{p}(\mathbf{x}_{k}^{\mathrm{3D}})$, defined by stacking the perspective projection of individual state variables as
$\mathbf{p}(\mathbf{x}_{k}^{\mathrm{3D}})=\frac{1}{\texttt{|px|}}\cdot\frac{f}{z_{k}}\left[\begin{smallmatrix}x_{k}\\ \left(\dot{x}_{k}-\dot{z}_{k}\cdot x_{k}\right)\\ y_{k}\\ \left(\dot{y}_{k}-\dot{z}_{k}\cdot y_{k}\right)\\ \omega_{k}\\ \dot{z}_{k}\cdot\omega_{k}\\ h_{k}\\ \dot{z}_{k}\cdot h_{k}\end{smallmatrix}\right]+\left[\begin{smallmatrix}[\mathbf{r}_{\mathrm{I|F}}]_{1}\\ 0\\ [\mathbf{r}_{\mathrm{I|F}}]_{2}\\ 0\\ 0\\ 0\\ 0\\ 0\end{smallmatrix}\right].$ (82)
@@
4.1.3 Physical Interpretation and Setup of Model Parameters
The parameters specified above are suitable for describing pedestrians observed with a camera positioned in parallel with the ground or reasonably elevated121212 For the case of a top-view, where the height $h_{k}$ can be assumed to coincide with the width $\omega_{k}$, both processes can be modeled with the mean $m_{\omega}=0.85$ m, standard deviation $\sigma_{\omega}=\tfrac{0.45}{3}$ m, and time-constant $\tau_{\omega}=0.4$. , i.e., if the height of the planar 3D BB state corresponds to the pedestrian’s height well. However, the physical modeling used in this paper allows for concise parameter selection, even for different situations.
[…]

### v2

@@ Abstract
A first-principle single-object model is proposed for pedestrian tracking. It is assumed that the extent of the moving object can be described via known statistics in 3D, such as pedestrian height. The proposed model thus need not constrain the object motion in 3D to a common ground plane, which is usual in 3D visual tracking applications. A nonlinear filter for this model is implemented using the unscented Kalman filter (UKF) and tested using the publicly available MOT-17 dataset. The proposed solution yields promising results in 3D while maintaining excellent results when projected into the 2D image. Moreover, the estimation error covariance matches the true one. Unlike conventional methods, the introduced model parameters have convenient meaning and can readily be adjusted for a problem.
Index Terms:
@@ I Introduction
This paper develops a linear stochastic dynamic model and the corresponding nonlinear measurement model suitable for the monocular pedestrian VTS, that (i) allows unconstrained object motion description in 3D space and (ii) does not rely on the common ground plane assumption. Instead, the proposed BB dynamics model is designed under the assumption of a pedestrian with average body proportions. This idea shares some similarities with [8, 14, 15]. In addition, techniques for setting and identifying the resulting state-space model parameters are discussed and illustrated using available datasets. A numerically stable unscented Kalman filter (UKF) is designed for the proposed model and is analyzed using the MOT-17 dataset [16, 17]. The results are illustrated in Fig. 1.
(a) Filtering results in 2D.
(b) Filtering results in 3D, $x$-$z$ view.
The paper is organized as follows. Section II provides background for monocular VTSs. Traditional tracking algorithms working in 2D are reviewed in Section III and the proposed method is presented in Section IV. A practical comparison is given in Section V, and the paper concludes in Section VI.
II Formalization of the Problem
This section outlines the relation between a moving object in 3D and its observation via a VDN applied to a camera image. The motion in 3D is discussed, followed by the projection of the object at a particular time into a pinhole camera image. Relation of the projection to the VDN output is described last.
II-A Object Motion in 3D
@@
II-B Pinhole Camera Model
@@
II-B1 Projection of a 3D Point into Focal Plane
@@
II-B2 Basis Change to Image Plane
@@
where $\mathbf{r}_{\mathrm{I},\mathrm{F}}$ expressed in pixels is the vector pointing from the origin of the image plane $\mathrm{I}$ to the center of the projection $\mathrm{F}$ as illustrated in Fig. 2b. The velocity $\dot{\mathbf{r}}_{\mathrm{F},\mathrm{P}}$, the length $s_{\mathrm{F}}$ and its temporal change rate $\dot{s}_{\mathrm{F}}$ are transformed from $\mathrm{F}$ to $\mathrm{I}$ by
@@
II-C Object’s Bounding Box and Annotation
@@
II-D Bounding Box Detection Modeling
@@
III Baseline Filtering Techniques in 2D
@@
III-A 2D-Model-Based Linear Time-Invariant Filter
Papers [5] and [6] present linear time-invariant motion222 The specific model used in this paper is referred to as BB dynamic model 2 in [5], and the notation used herein is slightly varied. and measurement models suitable for VTSs, including identification of their parameters. The motion model from [5] uses the nearly constant velocity (NCV) model, which is described first. Using it as a building block, the state-space model is introduced, followed by the construction of a filter for this model. Note that this resulting filter has not yet been presented elsewhere.
III-A1 Nearly Constant Velocity Model
@@
with the state noise covariance being a function of the sampling period $T=\nicefrac{{1}}{{\texttt{frame\_\,rate}}}$ in seconds and a parameter $q$ in m2s-3.
III-A2 State-Space Modeling Based on [5, 6]
@@
$\mathbf{Q}^{\mathrm{2D}}=\text{cov}\!\left(\,\mathbf{w}_{k}^{\mathrm{2D}}\,\right)=\gamma^{2}\cdot\text{diag}\left(\,{q_{\dot{\mathsf{x}}}\mathbf{T},q_{\dot{\mathsf{y}}}\mathbf{T},q_{\dot{\upomega}}\mathbf{T},q_{\dot{\mathsf{h}}}\mathbf{T}}\,\right),$ (15)
@@
III-A3 2D-Model-Based Filter for (12)
An optimal unbiased state estimator having minimal mean squared error (MSE) for the linear system such as (12) can be derived in a closed form. It is the Kalman filter (KF), formed by a recursion composed of the prediction and filtering steps.
@@
where $\mathbf{K}_{k}^{\mathrm{2D}}$ is the Kalman gain. The error corresponding to the predictive and filtering estimates $\mathbf{x}_{k|k\!-\!1}^{\mathrm{2D}}$ and $\mathbf{x}_{k|k}^{\mathrm{2D}}$, respectively, are characterized by the predictive and filtering covariance matrices $\mathbf{P}_{k|k\!-\!1}^{\mathrm{2D}}$ and $\mathbf{P}_{k|k}^{\mathrm{2D}}$, respectively. The filtering step (18) is performed only when a measurement $\mathbf{z}_{k}$ exists at the time step $k$, which generally applies throughout the paper.
III-A4 Filter Initialization
For a fair comparison, the filter is initialized using the measurement $\mathbf{z}_{0}$ as666 The matrix $\mathbf{H}$ appearing in (19) is numerically equal to $\mathbf{H}$ (14), but its even columns are in s-1, instead of s. Notation is not altered for simplicity.
@@ III-A4 Filter Initialization
  • The depth can be estimated777 The dept is estimated using (4) with $s_{\mathrm{O}}$ being the pedestrian height, substituted with the assumed mean pedestrian height $m_{h}$. Using (6), $s_{\mathrm{I}}$ is then replaced with the measured height. Note that the height is preferable to the width for the depth estimation, as the width of a pedestrian, including its hands, may change significantly due to hand swinging. as $z_{0|0}=\tfrac{f\cdot m_{h}}{\texttt{|px|}\cdot[\mathbf{z}_{0}]_{4}}$.
@@
III-B Bag-of-Tricks-Based Filter
@@
A notable adjustment was introduced in DeepSORT [21], which suggested that the covariance matrices $\mathbf{Q}$ and $\mathbf{R}$, appearing in the Kalman filter should be functions of some estimated state or measurement elements (see their code). Presumably, this should mimic the perspective projection effects. Small objects presumably in the background are given small $\mathbf{Q}$ and $\mathbf{R}$, and vice-versa. The resulting filter, however, is no longer a Kalman filter. The final adjustments introduced in BoT-SORT include changing the original state vector888 The original presentation from BoT-SORT paper [4] and the corresponding code was revisited, leading to alternative notation in this paper. to be
@@
thus all units of the state variables in $\mathbf{x}_{k}^{\mathrm{BoT}}$ (III-B) are pixels.
@@
$\mathbf{F}^{\mathrm{BoT}}$ $=\text{diag}\left(\,{\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right],\left[\begin{smallmatrix}1&1\\ 0&1\end{smallmatrix}\right]}\,\right).$
@@
Notice that the 2D-Model-Based Filter III-A completely neglects any 3D modeling, while the Bag-of-Tricks-Based Filter III-B takes it heuristically into account up to some level. None of these filters produce estimates in 3D. These issues are tackled in the following Section by modeling directly in 3D.
IV Proposed Model-Based Nonlinear Filter in 3D
@@ IV-A Geometric Modeling in 3D
The planar 3D BB is a perfectly aligned rectangle with the focal plane, as shown in Fig. 3. All the vertices and edges of the planar 3D BB have the same depth in front of the camera. $\square$
@@
where the position $x_{k}$, $y_{k}$, $z_{k}$ of the lower-bottom center of the box and its width $\omega_{k}$ and height $h_{k}$ are in meters, while the velocities $\dot{x}_{k}$, $\dot{y}_{k}$, $\dot{z}_{k}$ are in meters per second. An example of such a 3D BB state is given in Fig. 3, in which case position is $\mathbf{r}_{\mathrm{C,O}}=[x_{k}\ y_{k}\ z_{k}]^{\top}$, the velocity is $\dot{\mathbf{r}}_{\mathrm{C,O}}=[\dot{x}_{k}\ \dot{y}_{k}\ \dot{z}_{k}]^{\top}$, the height is $s_{\mathrm{O}}=h_{k}$, and the width is not denoted in Fig. 3.
The dynamics (1) of the 3D state is approximated as follows. In accordance with standard motion modeling of physical objects [1], an NCV model from Section III-A1 is used to describe the position/velocity of the state $\mathbf{x}_{k}^{\mathrm{3D}}$ (30). However, unlike the 2D case, the NCV model does not seem suitable for describing the extent. Therefore, this paper proposes to model the width and height using the following model.
IV-A1 Auto-Regressive Model
@@
where $\tau>0$ is a time-constant101010 Parameter $\tau$ can be specified as the elapsed time during which the process $\mathtt{p}(t)$, starting at $\mathtt{p}(0)$, changes its mean value to $\mathtt{p}(\tau)\approx 0.4\mathtt{p}(0)+0.6m$. , $w(t)$ is a zero-mean continuous white noise, and $\delta(t)$ is the Dirac delta function111111 Similarly to (9), a correct mathematical notation of (31) would involve a stochastic differential equation and a Wiener process [22]. .
@@
where $\alpha=e^{\nicefrac{{-\!T}}{{\tau}}}$. The above discrete model is usually called the auto-regressive (AR) model of order 1.
IV-A2 State-Space Modeling in 3D
@@
respectively, where $\mathbf{0}_{m\times n}$ is a zero matrix of dimensions $m\times n$, and where $\alpha_{\omega}=e^{\nicefrac{{-\!T}}{{\tau_{\omega}}}}$ and $\alpha_{h}=e^{\nicefrac{{-\!T}}{{\tau_{h}}}}$. The parameters of the width and height AR models are as follows. The height of the planar 3D BB is modeled as that of a pedestrian, using the mean $m_{h}=1.65$ m and time-constant $\tau_{h}\approx 4$ s. The width of the planar 3D BB is modeled with the mean $m_{\omega}=0.85$ m, and time-constant $\tau_{\omega}\approx 0.4$ s to account for rough changes due to the hand swinging. The matrix $\mathbf{Q}^{\mathrm{3D}}=\text{cov}\!\left(\,\!\mathbf{w}_{k}^{\mathrm{3D}}\!\,\right)$ is
@@
where the PSDs $q_{\dot{x}}\!=\!q_{\dot{y}}\!=\!q_{\dot{z}}\!=\!1$ m2s-3 are set for pedestrians  [23, pp. 418]. The standard deviations $\sigma_{\omega}\!=\!\tfrac{0.45}{3}$ m and $\sigma_{h}\!=\!\tfrac{0.3}{3}$ m of the width and height, respectively, are selected based on their assumed relation to the maximal deviations from the means (not velocities as in Section III-A4), i.e., $\sigma_{\omega}^{\mathrm{max}}=0.45$ m and $\sigma_{h}^{\mathrm{max}}=0.3$ m, respectively.
The nonlinear measurement equation (33b) is described by the function $\mathbf{x}_{k}^{\mathrm{2D}}=\mathbf{p}(\mathbf{x}_{k}^{\mathrm{3D}})$, defined by121212 The function is constructed by stacking the perspective projection of individual state variables. According to the AR model, the variables $\dot{\omega}_{k}$ and $\dot{h}_{k}$ are undefined and (4)-(6) is thus approximated with only $\dot{s}_{\mathrm{I}}=-\tfrac{f\dot{z}}{\texttt{|px|}z^{2}}s_{\mathrm{O}}$.
$\mathbf{p}(\mathbf{x}_{k}^{\mathrm{3D}})=\tfrac{f}{\texttt{|px|}\cdot z_{k}}\left[\begin{smallmatrix}x_{k}\\ \dot{x}_{k}-\dot{z}_{k}\cdot x_{k}/z_{k}\\ y_{k}\\ \dot{y}_{k}-\dot{z}_{k}\cdot y_{k}/z_{k}\\ \omega_{k}\\ -\dot{z}_{k}\cdot\omega_{k}/z_{k}\\ h_{k}\\ -\dot{z}_{k}\cdot h_{k}/z_{k}\end{smallmatrix}\right]+\left[\begin{smallmatrix}[\mathbf{r}_{\mathrm{I|F}}]_{1}\\ 0\\ [\mathbf{r}_{\mathrm{I|F}}]_{2}\\ 0\\ 0\\ 0\\ 0\\ 0\end{smallmatrix}\right].$ (37)
@@
IV-A3 Physical Interpretation and Setup of Model Parameters
The parameters specified above are suitable for describing pedestrians observed with a camera positioned in parallel with the ground or reasonably elevated131313 For the case of a top-view, where the height $h_{k}$ can be assumed to coincide with the width $\omega_{k}$, both processes can be modeled with the mean $m_{\omega}=0.85$ m, standard deviation $\sigma_{\omega}=\tfrac{0.45}{3}$ m, and time-constant $\tau_{\omega}=0.4$. , i.e., if the height of the planar 3D BB state corresponds to the pedestrian’s height well. However, the physical modeling used in this paper allows for concise parameter selection, even for different situations.
In [24], the PSDs $q_{\dot{x}},q_{\dot{y}},$ and $q_{\dot{z}}$ were optimized for tracking pedestrians, suggesting values from 0.7 up to 0.77 m2s-3. For vehicles, the PSDs can be selected as 10 m2s-3 [23, pp. 418]. For any case, the PSDs could also be specified according to the maximal acceleration $\ddot{r}^{\mathrm{max}}$ in $\cdots$-2 as [1, pp. 44]